A) A series circuit consists of a 0.01 µF capacitor, a 0.143 µF capacitor, and a 200 V battery. Find the charge on each capacitor. Answer in units of µC. B) Find the charge on the 0.01 µF capacitor if they are reconnected in parallel across the battery. Answer in units of µC. C)Find the charge on the 0.143 µF capacitor if they are reconnected in parallel across the battery. Answer in units of µC.
Answer #1

a)

net capacitance

1/ C = 1/ 0.01 + 1/ 0.143

C = 9.346*10^-3 uF

net charge through circuit

q = C V = 9.346*10^-9* 200 = 1.87 uC

charge on each capacitor = 1.87 uC

=====

b)

q(0.01 uF) = 0.01*10^-6* 200= 2 uC

======

c)

q(0.143) = 0.143* 10^-6 * 200 = 28.6 uC

======

Comment in case any doubt, will reply for sure.. Goodluck